A circle contains two tangents with equations X=5 and Y=7.What could be the coordinates of the centre?

Grade:12

2 Answers

PRIYANKA TATA ASKIITIAN
13 years ago

the centre of the circle will lie on the angular bisectors of the given pair of tangents.

the angular bisectors will be x-5 = ±(y-7)

=> y = x+2 and y = 12-x

=> (α,α+2) and (α,12-α) will be the centres where α is any parameter...

with further data to the problem we can get the exact centres to the required circle

priyanka tata

Souradeep Majumder
13 years ago

THE POSSIBLE COORDINATES ARE (5,7) AND (-5,-7)

 

 

APPROVE MY ANSWER.........